OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 4) and whose slope is 1/2.

OpenStudy (aroub):

y-y1=m(x-x1) y-4=1/2(x-2) y-4=x-2/2 => y=x/2-3